Laura’s Notebook: Manchester, 6/12

Screening day- Arrived at 8:30am, geared up to screen through the larvae. Grace worked the screening table, Olivia managed the buckets and whatnot, I counted and re-stocked larvae, and Katie took video and images. Finished at ~2:00pm. Things went very smoothly and we got a good routine down. Here are snap-shots of the data:

imageimage

After restocking, all buckets were fed 700mL algae cocktail of: 1/4 Tet, 1/4 609, 1/2 Chagra.

Images: using my phone & the new cell phone adapter, Katie imaged and observed activity. Images were taken using my phone. Here are her notes:

img_0565img_0566img_0567img_0568

I fixed all screened larvae with Lugols and stored in fridge for size imaging (under the scope with the scale) at a later date.

Collected new larvae, counted, sampled, and stocked:

  • HL-6 Ambient
  • HL-10 Low
  • K-6 Ambient
  • SN-6 Low B
  • NF-6 Low A

Other activities:

  • Arrival inspection:
    • All buckets very clear, likely have been without food for several hours.
    • It appears the dripper was left out of NF-10 Ambient B since Saturday. Air stone was still in there; no apparent mortalities.
    • Aeration was not sufficiently high on the west side of the larval table (mostly on K & HL groups).
    • Temperature in the whole system has dropped to ~16degC, possibly due to it being pretty cold outside…?
  • Rinsed all larval catchment buckets
  • Checked on the upwelling tank outside; it is holding at 17degC
  • Gigas:
    • Grace & Olivia cleaned, inspected for morts (none)
    • Grace measured and adjusted flow rate to set a 1.2L/min per side
    • I increased temp 1degC; temp was reading 15degC, so I set heater to 16degC
    • Installed aquarium pump in header tank, fed with 490mL Reeds Shellfish Diet, and set dosing pump to 57

from The Shell Game http://ift.tt/2sbcJoR
via IFTTT

Advertisements